Question: If F = 9/5C +32 make C the subject. I know the answer is C = 5/9 (F -32) but I need to follow the rationale for this through on a step by step (or line by line) basis please
Thanks Bob

Answer by stanbon(75887)

F = (9/5)C +32
subtract 32 from both sides
(9/5)C = F-32
---
Multiply both sides by (5/9)
C = (5/9)(F-32)
====================
Cheers,
Stan H.